Grade 12Algebra

Sir please help me as soon as possible !!!!
Find the value of (1 + 1/3.22 + 1/5.24 + 1/7.26 +...............infinity
a) ​​loge ​3
B) 1/2 ln 3
c) 1/2 loge 3 - 1
d) ln 3 - 1

Ranjeet Gupta
7 Years agoGrade 12

1 Answer

Ranjeet Gupta
7 Years ago
loge(1+x)= x - x^2/2 + x^3/3 - x^4/4.......       [eq 1] 
loge(1-x)= - x - x^2/2 - x^3/3 - x^4/4........       [eqn 2]  
Subtracting eq 2 from eq 1,
loge(l+x) /(1-x) = 2( x + x^3/3 + x^5/5 + x^7/7......)
And now we put x=1/2 from the question condition 
Then loge(3) = 2(1/2 + 1/3*2^3 + 1/5*2^5 + 1/7*2^7........)
                   = (1 + 1/3.2^2 + 1/5.2^4 + 1/7.2^6....)
            ......... Dividing by 2 on b.s
 Hence correct ans. a) loge (3)